OpenStudy (anonymous):

Choose the correct term: A positive discriminant means a quadratic equation has _______ ________ solutions. A negative discriminant means a quadratic equation has _______ ________ solutions. A discriminant of zero means a quadratic equation has ________ solutions

OpenStudy (anonymous):

2, none, 1

OpenStudy (anonymous):

1.real solutions 2. immaginary solutions

OpenStudy (anonymous):

think so..

OpenStudy (anonymous):

3. Real and equal solns

OpenStudy (anonymous):

m i rite...... for 1st two

OpenStudy (anonymous):

on account of the quadratic formula has a \[\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] in if the thing under the radical is positive you get two solutions, if it is zero you get just \[\frac{-b}{2a}\] and if it negative you have minus sign under the radical so not a real number

OpenStudy (anonymous):

technically..the first one shld be real and distinct roots

OpenStudy (anonymous):

go it my friend......

OpenStudy (anonymous):

yeah I got it but this question A negative discriminant means a quadratic equation has _______ ________ solutions. is still giving me trouble what are the two solutions to the question

myininaya (myininaya):

2 non-real solutions

OpenStudy (anonymous):

immaginary roots as we get \[\sqrt{-1}\]

OpenStudy (anonymous):

such type

OpenStudy (anonymous):

under root negative no.

OpenStudy (anonymous):

I get it now thank you

OpenStudy (anonymous):

most welcum